Under identical conditions gaseous CO2 and CCl4 are allowed to effuse through a pinhole. If the rate of effusion of the CO2 is 6.3x10-2 mol/s, what is the rate of effusion of the CCl4?
1
Expert's answer
2020-11-20T08:03:47-0500

According to Graham's law,

"\frac{R_1}{R_2}=\sqrt{\frac{M_2}{M_1}}" ,

where R are rates of effusion, and M are molar masses of the corresponding gases.


Therefore,

"\frac{R_{CO_2}}{R_{CCl_4}}=\sqrt{\frac{M_{CCl_4}}{M_{CO_2}}}"


"\frac{6.3\cdot10^{-2}mol/s}{R_{CCl_4}}=\sqrt{\frac{153.82g/mol}{44.01g/mol}}"


"R_{CCl_4}=3.4\cdot10^{-2}mol/s"


Answer: "3.4\cdot10^{-2}mol/s"